Solve the following recurrence relation clearly showing all work and each step of the proof and the reasoning thereof. Be sure to state the base case and assumption. List the the rules that apply to each step.
(10 points) Solve the recurrence relation: F(1) = -1; F(2) = 7; F(n) = -3F(n-2) - 4(n-1), n23 7.
Answer #1

The characteristic function for the recurrence relation is given as:

x2 = -3 - 4x

=> x2 + 4x + 3 = 0

The roots of this equation are -1 and -3.

This means that

F(n) = a * (-1)n + b * (-3)n

Now, F(1) = -1 => -a - 3b = -1 => a + 3b = 1

And, F(2) = 7 => a + 9b = 7

Solving these two equations, we get

a = -2 and b = 1

Then, F(n) = -2 * (-1)n + (-3)n